OpenStudy (anonymous):

Which phrase could be represented by this expression? k/5+ 8 A.eight more than five times a number B.eight more than a fifth of a number C.five times a number plus eight D.one-fifth of the sum of a number and eight

OpenStudy (anonymous):

Order of operations works well here for your expression \[\frac{ k }{ 5 } + 8\] k = some number

OpenStudy (anonymous):

Further more we see that the 5 is in the denominator of that number so we can say it is not the full number but a fifth of the number and since it does not number the number K the independent variable 8 will always be added to that number so using that reasoning you should now be able to find the correct phrase

OpenStudy (anonymous):

c..thank you

OpenStudy (anonymous):

Not C

OpenStudy (anonymous):

no d my bad

OpenStudy (anonymous):

C is saying \[5 \times k + 8\]

OpenStudy (anonymous):

Incorrect, D is saying \[\frac{ k }{ 5 } = k +8\]

OpenStudy (anonymous):

really it kind fit what you said

OpenStudy (anonymous):

Your equation is really \[K \times \frac{ 1 }{ 5 } + 8\]

OpenStudy (anonymous):

Sum implies the addition and actually the equals sign as well

OpenStudy (anonymous):

the only thing i could it is B

OpenStudy (anonymous):

i could think it is B

OpenStudy (anonymous):

Correct, because if you have one-fifth of a number then you have it over 5. So for example if i had one-half of a number it would be over 2. Since the 8 is going to be added onto whatever that result is it will be 8 more than that one-fifth